# FreDN: Spectral Disentanglement for Time Series Forecasting via Learnable Frequency Decomposition

### v1

@@ Abstract
Time series forecasting is essential in a wide range of real world applications. Recently, frequency-domain methods have attracted increasing interest for their ability to capture global dependencies. However, when applied to non-stationary time series, these methods encounter the spectral entanglement and the computational burden of complex-valued learning. The spectral entanglement refers to the overlap of trends, periodicities, and noise across the spectrum due to spectral leakage and the presence of non-stationarity. However, existing decompositions are not suited to resolving spectral entanglement. To address this, we propose the Frequency Decomposition Network (FreDN), which introduces a learnable Frequency Disentangler module to separate trend and periodic components directly in the frequency domain. Furthermore, we propose a theoretically supported ReIm Block to reduce the complexity of complex-valued operations while maintaining performance. We also re-examine the frequency-domain loss function and provide new theoretical insights into its effectiveness. Extensive experiments on seven long-term forecasting benchmarks demonstrate that FreDN outperforms state-of-the-art methods by up to 10%. Furthermore, compared with standard complex-valued architectures, our real-imaginary shared-parameter design reduces the parameter count and computational cost by at least 50%.
@@ Introduction
Time series forecasting is essential in energy, finance, and traffic management. Time-domain forecasting models continue to make notable progress in recent years. MLP-based architectures like TimeMixer (wang2024timemixer) and SOFTS (han2024softs) leverage multiscale mixing or centralized channel fusion to improve efficiency and scalability. At the same time, frequency-domain methods have gained traction for their ability to capture global dependencies and suppress noise. Models like FreTS (yi2023frequency) suggest that the frequency spectrum offers MLPs a more complete view of the signal, facilitating the learning of global patterns. Recent works such as FITS (xu2024fits) and FreDF (wang2025fredf) further demonstrate that frequency-based representations enhance forecasting accuracy.
@@
Frequency-domain techniques have become increasingly prevalent in time series forecasting, with the Fourier transform (oppenheim1999discrete) playing a central role in analyzing spectral structures. However, when applied to non-stationary real world time series, frequency-domain methods face two major difficulties: spectral entanglement that hampers effective decomposition, and the computational burden of learning with complex-valued representations. spectral entanglement is the overlap of trend, periodic, and noise components in the frequency domain, as illustrated in Figure 1 (d). This issue arises due to two main factors. First, the finite lookback window length $L$ introduces spectral leakage: the resolution of the Discrete Fourier Transform (DFT) is limited by $L$ according to the Nyquist theorem (oppenheim1999discrete). Any true frequency that does not align with the discrete basis causes energy to spread across all frequencies. Second, non-stationary trends, which do not exhibit periodicity, cannot be accurately represented by a small set of frequencies, as illustrated in Figure 1 (b).
The spectral entanglement undermines the effectiveness of conventional trend-seasonality decomposition methods (cleveland1990stl), which are typically employed to address non-stationarity in time series. For example, the moving average (wu2021autoformer; wang2024timemixer) can be viewed as applying a fixed low-pass filter in the frequency domain, but this approach is frequency-index-based and fails to resolve spectral entanglement, as shown in Figure 1 (e). Frequency-domain approaches such as Top-$K$ selection (wu2021autoformer; zhou2022fedformer) retain only dominant frequencies by amplitude, often discarding leaked or overlapping components (Figure 1 (f)). Moreover, frequency-domain learning involves complex-valued coefficients. Prior works such as FEDformer (zhou2022fedformer) and FreTS (yi2023frequency) apply specialized matrix operations to complex numbers, but these increase model complexity and hinder the integration of standard real-valued neural networks.
@@
Time series forecasting plays a crucial role in fields such as finance, energy, and climate science. Classical statistical models like ARMA, ARIMA, and VAR (box2015time; makridakis1997arma; watson1994vector; brockwell1991time) rely on linear assumptions and are limited in capturing nonlinear dependencies. Deep learning methods significantly improve modeling capacity, with RNNs (salinas2020deepar), CNNs (bai2018empirical; liu2022scinet), and Transformers (wu2021autoformer; zhou2021informer; nie2023patchtst; liu2024itransformer; wen2023transformers) demonstrating strong performance on complex temporal patterns. Recent efforts emphasize efficiency and fairness through compact designs (wang2024timemixer; han2024softs; lin2024sparsetsf; qiu2024tfb; qiu2025duet; qiu2025DBLoss). Frequency-domain modeling has gained attention for its ability to capture global periodic structures and suppress noise (zhang2025timeseriesanalysisfrequency). Approaches like FreTS, FITS, and FreDF (yi2023frequency; xu2024fits; wang2025fredf) operate directly on spectral representations, while models such as CoST and FiLM (woo2022cost; zhou2022film) incorporate frequency-domain operations into neural architectures. Broader spectral paradigms, such as Fourier Neural Operators and wavelet CNNs (li2021fourier; fang2024spiking), demonstrate strong modeling capacity.
@@
To handle non-stationarity, time-domain methods often decompose sequences into trend, seasonal, and noise components. Classical approaches include differencing (oppenheim1999discrete), STL (cleveland1990stl), SSA (golyandina2001analysis), and EMD (huang1998empirical). While effective, these techniques are tailored to time-domain processing and do not adapt well to frequency-domain tasks. Non-stationary sequences pose additional challenges in the frequency domain, where overlapping spectral energy from trend and periodic components causes spectral entanglement (zhang2025timeseriesanalysisfrequency). Heuristic solutions like Top-$K$ frequency selection (wu2021autoformer; zhou2022fedformer) are limited by spectral leakage and lack structural interpretability. Though localized spectral methods like STFT (oppenheim1999discrete) offer partial solutions, they are constrained by fixed resolution or high computational cost. Unlike our FreDN, which targets spectral entanglement via trend-seasonal separation, recent methods like FreDo (sun2022fredo), FAITH (li2025faith), and FreqMoE (liu2025freqmoe) either model frequencies holistically or decompose signals differently without resolving frequency overlap. To address these limitations, we propose FreDN, a learnable frequency-domain decomposition framework that disentangles trend and periodic components via adaptive frequency disentangler, while preserving compatibility with real-valued networks.
@@
Reversible Instance Normalization. We normalize the input using Reversible Instance Normalization (RevIN) (kim2021reversible), following the common practice in many models (nie2023patchtst; liu2024itransformer; han2024softs).
@@
Due to the spectral entanglement across all frequencies, existing methods such as moving average and Top-$K$ selection (zhou2022fedformer; wu2021autoformer) fail to function effectively in the frequency domain. To address this issue, we design a learnable Frequency Disentangler that adaptively separates the entire spectrum into smooth trend and seasonal components, effectively alleviating spectral entanglement.
@@
While trend patterns are naturally modeled in the time domain, periodic characteristics are more effectively captured in the complex plane. Existing methods such as FreTS (yi2023frequency) and FITS (xu2024fits) commonly adopt standard complex-valued linear layers to process frequency components. However, this approach requires specially designed complex matrix multiplications, introducing additional implementation complexity and computational overhead.
@@
Model Metric ETT (avg) Weather Electricity Traffic
@@
If we set $W_{i}=0$, the operation degenerates to $\tilde{Y}=W_{r}\tilde{X}$, which corresponds to the projection strategy used in the proposed ReIm Block. Instead of relying on full complex-valued multiplications, ReIm Block adopts shared real-valued weights, significantly simplifying implementation while retaining sufficient expressiveness. The following theorem formalizes its representational capacity:
Theorem 2 (Expressiveness of ReIm Block).
@@
Intuitively, the complex plane is a two-dimensional space, so any complex number can be represented as a linear combination of two orthogonal complex numbers (i.e., with a $90^{\circ}$ phase difference).
@@
Here, we provide a gradient-based analysis to examine why the frequency-domain MAE loss, first introduced in FreDF (wang2025fredf), may offer more efficient optimization behavior.
@@
FreDF (wang2025fredf) claims that ”by transforming the label sequence into this orthogonal frequency domain, the dependence from label autocorrelation could be effectively mitigated”. This statement is statistically inaccurate. As a linear transformation, DFT preserves the correlation structure of residuals, unless the residuals are strictly independent. This misunderstands the role of orthogonality: orthogonality enables frequency separation, not decorrelation.
@@
Specifically, from the Parseval’s theorem (oppenheim1999discrete), the energy of the residual is preserved under DFT, which means $\|\epsilon\|_{2}^{2}=\|\tilde{\epsilon}\|_{2}^{2}.$ This implies that time-MSE and frequency-MSE differ only by a constant factor in both loss magnitude and gradient.
@@
Structure Metric ETT (avg) Weather Electricity Traffic
@@
Datasets. To evaluate the performance of FreDN, we conduct extensive experiments on seven widely used long-term time series forecasting benchmarks: ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, and Traffic (wu2021autoformer; wang2025fredf; zhou2021informer). Detailed dataset descriptions are provided in  Appendix A.
@@
Baselines. We compare our model against a set of representative Linear-based or MLP-based baselines, including SOFTS (han2024softs), TimeMixer (wang2024timemixer), DLinear (zeng2023transformers). Two Frequency-based methods include FreTS (yi2023frequency), FreDF (wang2025fredf). We also consider Transformer-based methods including iTransformer (liu2024itransformer) and PatchTST (nie2023patchtst).
Implementation. The baseline models are reproduced using the official training scripts provided by FreDF (wang2025fredf). All models are optimized using the Adam optimizer (kingma2014adam). For FreDN, we use only the frequency-domain MAE loss without the time and frequency domains fused loss structure employed by FreDF (wang2025fredf). To ensure fairness, we follow the common practice of selecting hyperparameters based on the prediction length. For all models, the same set of hyperparameters is applied across all input lengths $\{96,192,336,512,720\}$ for a given prediction length. All experiments are implemented in PyTorch (paszke2019pytorch) and conducted on a single NVIDIA GeForce RTX 3090 GPU with 24GB memory. More experimental details are provided in  Appendix B.
@@
Table 5 reports long-term forecasting results on seven benchmarks averaged over $\tau\in\{96,192,336,720\}$. Each model uses default settings, and for every prediction length, the best result is selected across input lengths $L\in\{96,192,336,512,720\}$. FreDN achieves the best performance on most datasets. On average MSE, it surpasses FreDF (wang2025fredf) and FreTS (yi2023frequency) by 10% and 14%, and outperforms recent time-domain MLPs SOFTS (han2024softs) and TimeMixer (wang2024timemixer) by 6%. Although our best results with lookback window selecting outperform TimeMixer++ (wang2025timemixer) (as shown in the  Appendix C), we exclude it for fairness, since TimeMixer++ (wang2025timemixer) adopts a fixed $L=96$ and its official implementation is not publicly available. The hyperparameter sensitivity analysis are provided in  Appendix C. We also report the standard deviation of FreDN performance under five runs with different random seeds in  Appendix E, which exhibits that the performance is stable. Moreover, we further provide the visualization analysis in  Appendix F.
@@
The comparison of different decomposition methods is shown in Table 2. FreDN adopts our proposed Frequency Disentangler. MovDN applies a moving average with a fixed window size of 25, which follows the strategy used in TimeMixer (wang2024timemixer). TopKDN selects the frequency components with the largest $K$ amplitudes as the periodic part. Detailed configurations of MovDN and TopKDN are provided in  Appendix B.
In Figure 3, we present the decomposition of real world sequence. Figure 3 (b) shows the trend spectrum obtained using the Frequency Disentangler. The trend obtained by moving average in Figure 3(c) exhibits significant sidelobe leakage, characterized by a dominant main lobe accompanied by multiple high-energy sidelobes. This structure does not reflect the true spectral property of the trend component, but rather arises as a systematic artifact introduced by the decomposition method. This is because the finite moving window causes spectral leakage in the frequency domain, which is a direct consequence of the Nyquist theorem (oppenheim1999discrete). Specifically, the moving average decomposition is equivalent to a filter with the following frequency response:
@@
In Section ReIm Block, we provided a theoretical justification for the ReIm Block. To further demonstrate its practical efficiency, we compare the parameter count and average training time of the proposed ReIm Block with a Complex-Linear structure. The Complex-Linear structure replaces the linear projections in MLP with complex-valued linears, where the detailed architecture is described in  Appendix B. Specifically, to implement a Complex-Linear, it requires two sets of parameters for real and imaginary parts, and follows the complex matrix rule (golub2013matrix):
@@ Conclusion
Appendix Appendix A Datasets Description
@@
(zhou2021informer)  [^1] https://github.com/zhouhaoyi/ETDataset It contains 7 factors of electricity transformer from July 2016 to July 2018, and is collected from two different electric transformers labeled with 1 and 2. ETTh1 and ETTh2 are recorded every hour, and ETTm1 and ETTm2 are recorded every 15 minutes.
@@
(wu2021autoformer)  [^2] https://www.bgc-jena.mpg.de/wetter/ It collects 21 meteorological indicators, such as humidity and air temperature, from the Weather Station of the Max Planck Biogeochemistry Institute in Germany in 2020. The data sampling interval is every 10 minutes.
@@
(wu2021autoformer)  [^3] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 It contains electricity consumption of 321 clients for long-term forecasting. It is collected since 01/01/2011. The data sampling interval is every 15 minutes.
@@
(wu2021autoformer)  [^4] http://pems.dot.ca.gov It contains hourly traffic data from 862 car lanes for long-term forecasting settings. It is collected from January 2015 to December 2016 with a sampling interval of every 1 hour.
Table 4: Detailed dataset descriptions. Channels denotes the number of variables. Dataset Size is given as (Train / Validation / Test). Frequency indicates the sampling interval.
Dataset Channels Size Frequency
ETTh1 7 (8545, 2881, 2881) Hourly
ETTh2 7 (8545, 2881, 2881) Hourly
ETTm1 7 (34665, 11521, 11521) 15 min
ETTm2 7 (34665, 11521, 11521) 15 min
Weather 21 (36792, 5271, 10540) 10 min
ECL 321 (18317, 2633, 5261) Hourly
Traffic 862 (12185, 1757, 3509) Hourly
Appendix Appendix B Implement Details

### v2

@@ Abstract
Time series forecasting is essential in a wide range of real world applications. Recently, frequency-domain methods have attracted increasing interest for their ability to capture global dependencies. However, when applied to non-stationary time series, these methods encounter the spectral entanglement and the computational burden of complex-valued learning. The spectral entanglement refers to the overlap of trends, periodicities, and noise across the spectrum due to spectral leakage and the presence of non-stationarity. However, existing decompositions are not suited to resolving spectral entanglement. To address this, we propose the Frequency Decomposition Network (FreDN), which introduces a learnable Frequency Disentangler module to separate trend and periodic components directly in the frequency domain. Furthermore, we propose a theoretically supported ReIm Block to reduce the complexity of complex-valued operations while maintaining performance. We also re-examine the frequency-domain loss function and provide new theoretical insights into its effectiveness. Extensive experiments on seven long-term forecasting benchmarks demonstrate that FreDN outperforms state-of-the-art methods by up to 10%. Furthermore, compared with standard complex-valued architectures, our real-imaginary shared-parameter design reduces the parameter count and computational cost by at least 50%. We have made our code publicly available at https://github.com/An-z-d/FreDN
@@ Introduction
Time series forecasting is essential in energy, finance, and traffic management. Time-domain forecasting models continue to make notable progress in recent years. MLP-based architectures like TimeMixer (Wang et al. 2024) and SOFTS (Han et al. 2024) leverage multiscale mixing or centralized channel fusion to improve efficiency and scalability. At the same time, frequency-domain methods have gained traction for their ability to capture global dependencies and suppress noise. Models like FreTS (Yi et al. 2023) suggest that the frequency spectrum offers MLPs a more complete view of the signal, facilitating the learning of global patterns. Recent works such as FITS (Xu et al. 2024) and FreDF (Wang et al. 2025a) further demonstrate that frequency-based representations enhance forecasting accuracy.
@@
Frequency-domain techniques have become increasingly prevalent in time series forecasting, with the Fourier transform (Oppenheim 1999) playing a central role in analyzing spectral structures. However, when applied to non-stationary real world time series, frequency-domain methods face two major difficulties: spectral entanglement that hampers effective decomposition, and the computational burden of learning with complex-valued representations. Spectral entanglement is the overlap of trend, periodic, and noise components in the frequency domain, as illustrated in Figure 1 (d). This issue arises due to two main factors. First, the finite lookback window length $L$ introduces spectral leakage: the resolution of the Discrete Fourier Transform (DFT) is limited by $L$ according to the Nyquist theorem (Oppenheim 1999). Any true frequency that does not align with the discrete basis causes energy to spread across all frequencies. Second, non-stationary trends, which do not exhibit periodicity, cannot be accurately represented by a small set of frequencies, as illustrated in Figure 1 (b).
The spectral entanglement undermines the effectiveness of conventional trend-seasonality decomposition methods (Cleveland et al. 1990), which are typically employed to address non-stationarity in time series. For example, the moving average (Wu et al. 2021; Wang et al. 2024) can be viewed as applying a fixed low-pass filter in the frequency domain, but this approach is frequency-index-based and fails to resolve spectral entanglement, as shown in Figure 1 (e). Frequency-domain approaches such as Top-$K$ selection (Wu et al. 2021; Zhou et al. 2022b) retain only dominant frequencies by amplitude, often discarding leaked or overlapping components (Figure 1 (f)). Moreover, frequency-domain learning involves complex-valued coefficients. Prior works such as FEDformer (Zhou et al. 2022b) and FreTS (Yi et al. 2023) apply specialized matrix operations to complex numbers, but these increase model complexity and hinder the integration of standard real-valued neural networks.
@@
Time series forecasting plays a crucial role in fields such as finance, energy, and climate science. Classical statistical models like ARMA, ARIMA, and VAR (Box et al. 2015; Makridakis and Hibon 1997; Watson 1994; Brockwell 1991) rely on linear assumptions and are limited in capturing nonlinear dependencies. Deep learning methods significantly improve modeling capacity, with RNNs (Salinas et al. 2020), CNNs (Bai et al. 2018; Liu et al. 2022), Transformers (Wu et al. 2021; Zhou et al. 2021; Nie et al. 2023; Liu et al. 2024; Wen et al. 2023), and other task-specific architectures (Du et al. 2020, 2021, 2023a, 2023b) demonstrating strong performance on complex temporal patterns. Recent efforts emphasize efficiency and fairness through compact designs (Wang et al. 2024; Han et al. 2024; Lin et al. 2024; Qiu et al. 2024, 2025b, 2025a). Frequency-domain modeling has gained attention for its ability to capture global periodic structures and suppress noise (Zhang et al. 2025). Approaches like FreTS, FITS, and FreDF (Yi et al. 2023; Xu et al. 2024; Wang et al. 2025a) operate directly on spectral representations, while models such as CoST and FiLM (Woo et al. 2022; Zhou et al. 2022a) incorporate frequency-domain operations into neural architectures. Broader spectral paradigms, such as Fourier Neural Operators and wavelet CNNs (Li et al. 2021; Fang et al. 2024), demonstrate strong modeling capacity.
@@
To handle non-stationarity, time-domain methods often decompose sequences into trend, seasonal, and noise components. Classical approaches include differencing (Oppenheim 1999), STL (Cleveland et al. 1990), SSA (Golyandina et al. 2001), and EMD (Huang et al. 1998). While effective, these techniques are tailored to time-domain processing and do not adapt well to frequency-domain tasks. Non-stationary sequences pose additional challenges in the frequency domain, where overlapping spectral energy from trend and periodic components causes spectral entanglement (Zhang et al. 2025). Heuristic solutions like Top-$K$ frequency selection (Wu et al. 2021; Zhou et al. 2022b) are limited by spectral leakage and lack structural interpretability. Though localized spectral methods like STFT (Oppenheim 1999) offer partial solutions, they are constrained by fixed resolution or high computational cost. Unlike our FreDN, which targets spectral entanglement via trend-seasonal separation, recent methods like FreDo (Sun and Boning 2022), FAITH (Li et al. 2025), and FreqMoE (Liu 2025) either model frequencies holistically or decompose signals differently without resolving frequency overlap. To address these limitations, we propose FreDN, a learnable frequency-domain decomposition framework that disentangles trend and periodic components via adaptive frequency disentangler, while preserving compatibility with real-valued networks.
@@
Reversible Instance Normalization. We normalize the input using Reversible Instance Normalization (RevIN) (Kim et al. 2021), following the common practice in many models (Nie et al. 2023; Liu et al. 2024; Han et al. 2024).
@@
Due to the spectral entanglement across all frequencies, existing methods such as moving average and Top-$K$ selection (Zhou et al. 2022b; Wu et al. 2021) fail to function effectively in the frequency domain. To address this issue, we design a learnable Frequency Disentangler that adaptively separates the entire spectrum into smooth trend and seasonal components, effectively alleviating spectral entanglement.
@@
Existing methods such as FreTS (Yi et al. 2023) and FITS (Xu et al. 2024) commonly adopt standard complex-valued linear layers to process frequency components. However, this approach requires specially designed complex matrix multiplications, introducing additional implementation complexity and computational overhead.
@@
Model Metric ETT Weather Electricity Traffic
@@
If we set $W_{i}=0$, the operation degenerates to $\tilde{Y}=W_{r}\tilde{X}$, which corresponds to the projection strategy used in the proposed ReIm Block. Instead of relying on full complex-valued multiplications, ReIm Block adopts shared real-valued weights, significantly simplifying implementation while retaining sufficiency. The following theorem formalizes its representational capacity:
Theorem 2 (ReIm Block).
@@
Intuitively, the complex plane is a two-dimensional space, so any complex number can be represented as a linear combination of two orthogonal complex numbers (i.e., with a $90^{\circ}$ phase difference). This shows ReIm retains ComplexLinear’s key feature: frequency interaction.
@@
Here, we provide a gradient-based analysis to examine why the frequency-domain MAE loss, first introduced in FreDF (Wang et al. 2025a), may offer more efficient optimization behavior.
@@
FreDF (Wang et al. 2025a) claims that ”by transforming the label sequence into this orthogonal frequency domain, the dependence from label autocorrelation could be effectively mitigated”. This statement is statistically inaccurate. As a linear transformation, DFT preserves the correlation structure of residuals, unless the residuals are strictly independent. This misunderstands the role of orthogonality: orthogonality enables frequency separation, not decorrelation.
@@
Specifically, from the Parseval’s theorem (Oppenheim 1999), the energy of the residual is preserved under DFT, which means $\|\epsilon\|_{2}^{2}=\|\tilde{\epsilon}\|_{2}^{2}.$ This implies that time-MSE and frequency-MSE differ only by a constant factor in both loss magnitude and gradient.
@@
Structure Metric ETT Weather Electricity Traffic
@@
Datasets. To evaluate the performance of FreDN, we conduct extensive experiments on seven widely used long-term time series forecasting benchmarks: ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, and Traffic (Wu et al. 2021; Wang et al. 2025a; Zhou et al. 2021). Detailed dataset descriptions are provided in Appendix A.
@@
Baselines. We compare our model against a set of representative Linear-based or MLP-based baselines, including SOFTS (Han et al. 2024), TimeMixer (Wang et al. 2024), DLinear (Zeng et al. 2023). Two Frequency-based methods include FreTS (Yi et al. 2023), FreDF (Wang et al. 2025a). We also consider Transformer-based methods including iTransformer (Liu et al. 2024) and PatchTST (Nie et al. 2023).
Implementation. The baseline models are reproduced using the official training scripts provided by FreDF (Wang et al. 2025a). All models are optimized using the Adam optimizer (Kingma and Ba 2015). For FreDN, we use only the frequency-domain MAE loss without the time and frequency domains fused loss structure employed by FreDF (Wang et al. 2025a). To ensure fairness, we follow the common practice of selecting hyperparameters based on the prediction length. For all models, the same set of hyperparameters is applied across all input lengths $\{96,192,336,512,720\}$ for a given prediction length. All experiments are implemented in PyTorch (Paszke et al. 2019) and conducted on a single NVIDIA GeForce RTX 3090 GPU with 24GB memory. More experimental details are provided in Appendix B.
@@
Table 1 reports long-term forecasting results on seven benchmarks averaged over $\tau\in\{96,192,336,720\}$. Each model uses default settings, and for every prediction length, the best result is selected across input lengths $L\in\{96,192,336,512,720\}$. FreDN achieves the best performance on most datasets. On average MSE, it surpasses FreDF (Wang et al. 2025a) and FreTS (Yi et al. 2023) by 10% and 14%, and outperforms recent time-domain MLPs SOFTS (Han et al. 2024) and TimeMixer (Wang et al. 2024) by 6%. Although our best results with lookback window selecting outperform TimeMixer++ (Wang et al. 2025b) (as shown in the Appendix C), we exclude it for fairness, since TimeMixer++ (Wang et al. 2025b) adopts a fixed $L=96$ and its official implementation is not publicly available. The hyperparameter sensitivity analysis are provided in Appendix C. We also report the standard deviation of FreDN performance under five runs with different random seeds in Appendix E, which exhibits that the performance is stable. We further provide the visualization analysis in Appendix F.
@@
The comparison of different decomposition methods is shown in Table 2. FreDN adopts our proposed Frequency Disentangler. MovDN applies a moving average with a fixed window size of 25, which follows the strategy used in TimeMixer (Wang et al. 2024). TopKDN selects the frequency components with the largest $K$ amplitudes as the periodic part. Detailed configurations of MovDN and TopKDN are provided in Appendix B.
In Figure 3, we present the decomposition of real world sequence. Figure 3 (b) shows the trend spectrum obtained using the Frequency Disentangler. The trend obtained by moving average in Figure 3(c) exhibits significant sidelobe leakage, characterized by a dominant main lobe accompanied by multiple high-energy sidelobes. This structure does not reflect the true spectral property of the trend component, but rather arises as a systematic artifact introduced by the decomposition method. This is because the finite moving window causes spectral leakage in the frequency domain, which is a direct consequence of the Nyquist theorem (Oppenheim 1999). Specifically, the moving average decomposition is equivalent to a filter with the following frequency response:
@@
In Section ReIm Block, we provided a theoretical justification for the ReIm Block. To further demonstrate its practical efficiency, we compare the parameter count and average training time of the proposed ReIm Block with a Complex-Linear structure. The Complex-Linear structure replaces the linear projections in MLP with complex-valued linears, where the detailed architecture is described in Appendix B. Specifically, to implement a Complex-Linear, it requires two sets of parameters for real and imaginary parts, and follows the complex matrix rule (Golub and Van Loan 2013):
@@ Conclusion
Appendix A Datasets Description
@@
(Zhou et al. 2021)  [^1] https://github.com/zhouhaoyi/ETDataset It contains 7 factors of electricity transformer from July 2016 to July 2018, and is collected from two different electric transformers labeled with 1 and 2. ETTh1 and ETTh2 are recorded every hour, and ETTm1 and ETTm2 are recorded every 15 minutes.
@@
(Wu et al. 2021)  [^2] https://www.bgc-jena.mpg.de/wetter/ It collects 21 meteorological indicators, such as humidity and air temperature, from the Weather Station of the Max Planck Biogeochemistry Institute in Germany in 2020. The data sampling interval is every 10 minutes.
@@
(Wu et al. 2021)  [^3] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 It contains electricity consumption of 321 clients for long-term forecasting. It is collected since 01/01/2011. The data sampling interval is every 15 minutes.
@@
(Wu et al. 2021)  [^4] http://pems.dot.ca.gov It contains hourly traffic data from 862 car lanes for long-term forecasting settings. It is collected from January 2015 to December 2016 with a sampling interval of every 1 hour.
Table 4: Detailed dataset descriptions. Channels denotes the variate number in each dataset. Dataset Size denotes the total number of time points in (Train, Validation, Test) split respectively. Prediction Horizon denotes the forcasting length and four prediction settings are included in each dataset. Frequency denotes the sampling interval of time points.
Dataset Channels Prediction Horizon Dataset Size Frequency
ETTh1 7 96, 192, 336, 720 (8545, 2881, 2881) Hourly
ETTh2 7 96, 192, 336, 720 (8545, 2881, 2881) Hourly
ETTm1 7 96, 192, 336, 720 (34665, 11521, 11521) 15min
ETTm2 7 96, 192, 336, 720 (34665, 11521, 11521) 15min
Weather 21 96, 192, 336, 720 (36792, 5271, 10540) 10min
ECL 321 96, 192, 336, 720 (18317, 2633, 5261) Hourly
Traffic 862 96, 192, 336, 720 (12185, 1757, 3509) Hourly
Appendix B Implement Details